A certain brand of razor blades c omes in packages of 6, 12, and 24 blades, costing $2, $3, $4 per packages. a store sold 12 packages containing a total of 162 blades razor blades and took in $35. How many packages of each types of each types were sold?
Answer by stanbon(75887)

A certain brand of razor blades c omes in packages of 6, 12, and 24 blades, costing $2, $3, $4 per packages. a store sold 12 packages containing a total of 162 blades razor blades and took in $35. How many packages of each types of each types were sold?
-----------------
Let "x" be # of 6, "y" the # of 12, and "z" be # of 24 blade packages.
------------------------------
x+y+z = 12 packages
6x+12y+24z = 162 blades
2x+3y+4z = 35
=====================================
Using the TI-83 Matrix function I find:
x = 5
y = 3
z = 4
===========================
Cheers,
Stan H.